### v2

# Additive actions on projective hypersurfaces with a finite number of orbits

Abstract.
An induced additive action on a projective variety $X\subseteq\mathbb{P}^{n}$ is a regular action of the group $\mathbb{G}_{a}^{m}$ on $X$ with an open orbit, which can be extended to a regular action on the ambient projective space $\mathbb{P}^{n}$. In this work, we classify all projective hypersurfaces admitting an induced additive action with a finite number of orbits.
Key words and phrases:
Algebraic variety, algebraic group, unipotent group, algebraic group actions, open orbit, additive action, local algebra, projective hypersurface
2020 Mathematics Subject Classification:
Primary 14L30, 14J70; Secondary 13E10.
The work was supported by the grant RSF 23-71-01100
1. Introduction
Let ${\mathbb{K}}$ be an algebraically closed field of zero characteristic. By a variety or an algebraic group we always mean an algebraic variety or an algebraic group over ${\mathbb{K}}$. By open and closed subsets of algebraic varieties we always mean open and closed subsets in Zariski topology. We denote by ${\mathbb{G}}_{a}=({\mathbb{K}},+)$ the additive group of the ground field and by ${\mathbb{G}}_{a}^{m}$ the group
${\mathbb{G}}_{a}^{m}=\underbrace{{\mathbb{G}}_{a}\times\cdots\times{\mathbb{G}}_{a}}_{\text{$m$ times}}.$
Definition 1.
An additive action on an algebraic variety $X$ is a regular effective action of ${\mathbb{G}}_{a}^{m}$ on $X$ with an open orbit. By an induced additive action on an embedded projective algebraic variety $X\subseteq{\mathbb{P}}^{n}$ we mean a regular effective action of ${\mathbb{G}}_{a}^{m}$ on ${\mathbb{P}}^{n}$ such that the variety $X$ is the closure of an orbit of ${\mathbb{G}}_{a}^{m}$.
Not every additive action on a projective variety is induced. An example can be found in [4, Example 1]. However, when the projective variety $X\subseteq{\mathbb{P}}^{n}$ is normal and linearly normal, then every additive action of ${\mathbb{G}}_{a}^{m}$ on $X$ lifts to the regular effective action of ${\mathbb{G}}_{a}^{m}$ on the projective space ${\mathbb{P}}^{n}$.
In [19] a remarkable correspondence between additive actions on the projective space ${\mathbb{P}}^{n}$ and local algebras of dimension $n+1$ was obtained. By a local algebra we mean a commutative associative algebra over ${\mathbb{K}}$ with a unit and a unique maximal ideal. We will recall this correspondence in Section 2. A correspondence between actions of arbitrary commutative algebraic groups on ${\mathbb{P}}^{n}$ with an open orbit and associative commutative algebras with a unit element of dimension $n+1$ was established in [20].
The systematic study of additive actions on projective and complete varieties was initiated in [23, 6, 4]. There are several results on additive actions on projective hypersurfaces. For example, it was proven in [23] that there is a unique additive action on a non-degenerate quadric. This result was generalized in [11], where actions of arbitrary algebraic commutative groups on non-degenerate quadrics with an open orbit were described. In [4] and [6] induced actions on projective hypersurfaces were studied. It was proven in [7] that a non-degenerate hypersurface (see Definition 2) admits at most one additive action. When a degenerate hypersurface admits an additive action, then there are at least two non-isomorphic additive actions on it, see [10]. For additive actions on degenerate hypersurfaces we refer also to [21].
Flag varieties admitting an additive action were classified in [1] and all additive actions on flag varieties were classified in [16]. Additive actions on toric varieties were studied in [5, 3, 17, 18, 22, 24, 25]. There are results on additive actions on Fano varieties in [14, 15, 8, 9, 26]. For a detailed review of the results on additive actions we refer to [7].
Among actions of algebraic groups on algebraic varieties, actions with a finite number of orbits are of particular interest. For example, toric varieties can be characterized as varieties on which an algebraic torus acts with a finite number of orbits. Spherical varieties admit an action of a reductive group with a finite number of orbits. Additive actions with a finite number of orbits on complete varieties, with an additional condition on the actions of one-dimensional subgroups, were described in the work [13]; see also Section 4.4.
[…]
Our final results are stated in Theorem 3 and Corollary 4. Geometrically they mean the following.
  • a)
    There is exactly one curve in ${\mathbb{P}}^{2}$ which admits an induced additive action with a finite number of orbits. It is ${\mathbb{P}}^{1}$ embedded in ${\mathbb{P}}^{2}$ via Veronese embedding of degree 2.
  • b)
    There are exactly three surfaces in ${\mathbb{P}}^{3}$ which admit an induced additive action with a finite number of orbits. They are
    1. (1)
      ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ embedded in ${\mathbb{P}}^{3}$ via Segre embedding;
    2. (2)
      the non-degenerate cubic;
    3. (3)
      the degenerate hypersurface of degree 2 which is the projective cone over the hypersurface from the point a).
  • c)
    When $n>3$, there are exactly two hypersurfaces in ${\mathbb{P}}^{n}$ which admit an induced additive action with a finite number of orbits. One of them is a non-degenerate hypersurface $X_{n}$ of degree $n$. The other one is a degenerate hypersurface $Y_{n-1}$ of degree $n-1$ which is the projective cone over $X_{n-1}.$
The structure of the text is as follows. In Section 2 we recall known facts about additive actions on projective space and projective hypersurfaces. In Section 3 we prove the main results and find all projective hypersurfaces that admit an additive action with a finite number of orbits. Finally, in Section 4 we discuss properties of obtained hypersurfaces such as the structure of orbits, smoothness, normality and the total number of different additive actions.
Acknowledgments.
The authors are grateful to the anonymous reviewer for careful reading and valuable suggestions for improving the text.
2. Additive actions on projective varieties
In this section we recall some facts on additive actions on projective varieties. We say that two induced additive actions on a projective variety $X\subseteq{\mathbb{P}}^{n}$ are equivalent if one is obtained from the other via an automorphism of ${\mathbb{P}}^{n}$ preserving $X$.
Proposition 1.
[19, Proposition 2.15] There is a one-to-one correspondence between
  1. (1)
    equivalence classes of additive actions on ${\mathbb{P}}^{n}$;
  2. (2)
    isomorphism classes of local algebras of dimension $n+1$.
[…]
$m\mapsto\exp(m)=\sum_{i\geq 0}\frac{m^{i}}{i!},\;\text{ for }m\in\mathfrak{m}.$
This map is well-defined on $\mathfrak{m}$. The additive group of $\mathfrak{m}$ is isomorphic to ${\mathbb{G}}_{a}^{n}$ and $\mathfrak{m}$ acts on the algebra $A$ by the following rule: $m\circ a=\exp(m)\cdot a$. This is an algebraic action. The stabilizer of a unit is trivial, so we have the following isomorphisms of algebraic varieties
${\mathbb{A}}^{n}\simeq{\mathbb{G}}_{a}^{n}\simeq\mathrm{exp}(\mathfrak{m})\cdot 1=1+\mathfrak{m},$
where the last equality is satisfied since the map
$1+m\mapsto\ln(1+m)=\sum_{i>0}(-1)^{i-1}\frac{m^{i}}{i},\;\text{ for }m\in\mathfrak{m},$
is well-defined on $1+\mathfrak{m}$ and $\exp(\ln(1+m))=1+m$. The action of $\mathfrak{m}$ on the algebra $A$ defines an algebraic action of ${\mathbb{G}}_{a}^{n}$ on the projective space ${\mathbb{P}}^{n}={\mathbb{P}}(A)$ by the rule
$m\circ p(a)=p(\exp(m)\cdot a),$
where the map $p\colon A\setminus\{0\}\to{\mathbb{P}}(A)$ is the canonical projection. The orbit of $p(1)$ is the open orbit, so this defines an additive action on ${\mathbb{P}}^{n}$. See [7, Example 1.50] for further examples of this construction.
Proposition 2.
[4, Proposition 3] There is a one-to-one correspondence between
  1. (1)
    equivalence classes of pairs $(X,\alpha),$ where $X$ is a hypersurface in $\mathbb{P}^{n}$ and $\alpha$ is an induced additive action on $X$;
  2. (2)
[…]
The pairs $(A,U)$ from Proposition 2 are called $H$-pairs. We say that two $H$-pairs $(A_{1},U_{1})$ and $(A_{2},U_{2})$ are isomorphic if there is an isomorphism of local algebras $\varphi\colon A_{1}\to A_{2}$ such that $\varphi(U_{1})=U_{2}.$
[…]
$X=p(\overline{\exp(U)\cdot 1}).$
Then $X$ is a hypersurface in ${\mathbb{P}}(A)={\mathbb{P}}^{n}$ and the group $U$ acts on $X$ with an open orbit.
The following results illustrate how to find the defining equation and degree of $X$.
Theorem 1.
[6, Theorem 5.1] The degree of the hypersurface $X$ is equal to the largest number $d\in\mathbb{N}$ such that $\mathfrak{m}^{d}\nsubseteq U$, where $\mathfrak{m}$ is the maximal ideal in the corresponding local algebra $A$.
Theorem 2.
[7, Theorem 2.14] The hypersurface $X$ is given in ${\mathbb{P}}(A)$ by the following homogeneous equation:
$z_{0}^{d}\pi\left(\ln\left(1+\frac{z}{z_{0}}\right)\right)=0,$
where $z_{0}\in{\mathbb{K}},\,z\in\mathfrak{m}$ and $\pi\colon\mathfrak{m}\to\mathfrak{m}/U\simeq{\mathbb{K}}$ is the canonical projection.
It is also possible to describe elements $a\in A$ such that $\pi(a)\in X.$
Proposition 3.
[7, Corollary 2.18] The complement of the open $U$-orbit in $X$ is the set
$\{\,p(z)\mid z\in\mathfrak{m}\text{ such that }z^{d}\in U\,\},$
where $p\colon A\setminus\{0\}\to{\mathbb{P}}(A)$ is the canonical projection and $d$ is the degree of $X$.
Corollary 1.
Suppose that the point $x\in X$ belongs to the complement of the open orbit of the group $U$. Then the $\mathfrak{m}$-orbit of $x$ is contained in $X$.
Proof.
Let us take $z\in\mathfrak{m}$ such that $p(z)=x$ lies in $X$. Then $z^{d}\in\mathfrak{m}^{d}\cap U$. The $\mathfrak{m}$-orbit of the element $z$ is $z+z\cdot\mathfrak{m}$. But then $(z+z\cdot\mathfrak{m})^{d}\subseteq z^{d}+\mathfrak{m}^{d+1}\subseteq U$. So $p(z+z\cdot\mathfrak{m})\subseteq X$. ∎
We recall that a socle of a local algebra $A$ is the ideal $\mathrm{Soc}(A):=\{z\in A\mid z\cdot\mathfrak{m}=0\}$.
Corollary 2.
The set $\{\,p(z)\mid z\in\mathrm{Soc}(A)\setminus\{0\}\,\}$ is contained in $X$.
Proof.
For all $z\in\mathrm{Soc}(A)$ we have $z^{d}=0$ is in the group $U$. ∎
Corollary 3.
If $\mathrm{dim}(\mathrm{Soc}(A))>1$ then there are infinitely many $U$-orbits on $X$.
Proof.
If $z\in\mathrm{Soc}(A)$ then $\exp(U)\cdot z=\{z\}$. So the set $\{\,p(z)\mid z\in\mathrm{Soc}(A)\setminus\{0\}\,\}\subseteq X$ consists of the $U$-fixed points and has dimension at least 1. ∎
It is also possible to describe the relationship between $\mathfrak{m}$-orbits and $U$-orbits on $X$ For an element $z\in A$ we denote by $\mathrm{Ann}(z)$ the ideal $\{a\in A\mid az=0\}.$
Proposition 4.
Let $z\in\mathfrak{m}\setminus\{0\}$ be an element with $p(z)\in X$.
  1. (1)
    If $\mathrm{Ann}(z)+U=\mathfrak{m}$, then the $\mathfrak{m}$-orbit of $p(z)$ coincides with the $U$-orbit.
  2. (2)
    Otherwise, $\mathrm{Ann}(z)\subseteq U$ and the $\mathfrak{m}$-orbit of $p(z)$ is the union of an infinite number of $U$-orbits.
Proof.
We will show that $\mathrm{Ann}(z)$ coincides with the stabilizer $\mathrm{St}_{\mathfrak{m}}(p(z))$ with respect to the $\mathfrak{m}$-action. The inclusion $\mathrm{Ann}(z)\subseteq\mathrm{St}_{\mathfrak{m}}(p(z))$ is clear. Indeed, if $a\in\mathrm{Ann}(z)$ then we have $az=0$ and $\exp(a)\cdot z=z$. We now show the reverse inclusion. If $a\in\mathrm{St}_{\mathfrak{m}}(p(z))$ then
$\exp(a)\cdot p(z)=p(z+az+\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots)=p(z),$
which implies
$z+az+\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots=\lambda z.$
for some $\lambda\in{\mathbb{K}}.$ There is a number $k\in\mathbb{N}$ such that $z\in\mathfrak{m}^{k}\setminus\mathfrak{m}^{k+1}$. Then the element $az+\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots$ lies in the ideal $\mathfrak{m}^{k+1}$. So we have
$az+\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots\ =(\lambda-1)z,$
which is possible only when $\lambda=1.$ Therefore,
$az+\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots\ =0.$
If $az\neq 0$ then there is $r$ such that $az\in\mathfrak{m}^{r}\setminus\mathfrak{m}^{r+1}.$ But $\frac{a^{2}}{2}z+\frac{a^{3}}{6}z+\ldots\ \in\mathfrak{m}^{r+1}.$ So $az=0$ and $a\in\mathrm{Ann}(z).$
Thus, the $\mathfrak{m}$-orbit of $p(z)$ is isomorphic to $\mathfrak{m}/\mathrm{Ann}(z)$ and $U$-orbit of $p(z)$ is isomorphic to
$U/(\mathrm{Ann}(z)\cap U)\simeq(U+\mathrm{Ann}(z))/\mathrm{Ann}(z).$
Hence, if $U+\mathrm{Ann}(z)=\mathfrak{m}$ then the $\mathfrak{m}$-orbit of $p(z)$ coincides with the $U$-orbit, and if $U+\mathrm{Ann}(z)\neq\mathfrak{m}$ then the action of $U$ on $\mathfrak{m}/\mathrm{Ann}(z)$ has infinitely many orbits. Since the codimension of $U$ in $\mathfrak{m}$ is 1, in the last case we have $\mathrm{Ann}(z)\subseteq U.$
3. Main result
In this section we state our main result. Recall that for a local algebra $A$ with the maximal ideal $\mathfrak{m}$ the following sequence of numbers
$(\dim_{{\mathbb{K}}}A/\mathfrak{m},\ \dim_{{\mathbb{K}}}\mathfrak{m}/\mathfrak{m}^{2},\ \dim_{{\mathbb{K}}}\mathfrak{m}^{2}/\mathfrak{m}^{3},\ \ldots\ )$
is called a Hilbert-Samuel sequence.
Proposition 5.
Let $(A,U)$ be an $H$-pair and $X$ be the corresponding hypersurface in ${\mathbb{P}}(A).$ Suppose that there are finitely many $U$-orbits in $X$. Then the Hilbert-Samuel sequence of $A$ is either (1,1,1,,1)(1,1,1$\ldots$,1)( 1 , 1 , 1 , … , 1 ) or (1,2,1,1)(1,2,1$\ldots$ 1)( 1 , 2 , 1 , … 1 ).
Proof.
Since ${\mathbb{K}}$ is algebraically closed and $A/\mathfrak{m}$ is a finite-dimensional field over ${\mathbb{K}}$ we have
$\dim_{{\mathbb{K}}}A/\mathfrak{m}=1.$
Suppose there is a number $k\geq 2$ with $\dim_{{\mathbb{K}}}\mathfrak{m}^{k}/\mathfrak{m}^{k+1}>1.$ For all $z\in\mathfrak{m}^{k}\setminus\{0\}$ we have
$z^{d}\in\mathfrak{m}^{kd}\subseteq\mathfrak{m}^{d+1}\subseteq U,$
where $d$ is the degree of $X$. Then $p(z)$ lies in $X$ for all $z\in\mathfrak{m}^{k}\setminus\{0\}$. The $\mathfrak{m}$-orbit of $p(z)$ is $p(z+z\cdot\mathfrak{m})\subseteq p(z+\mathfrak{m}^{k+1})$. Thus, if the images of elements $z_{1}$ and $z_{2}$ from $\mathfrak{m}^{k}$ in $\mathfrak{m}^{k}/\mathfrak{m}^{k+1}$ are not proportional, then the $\mathfrak{m}$-orbits of $p(z_{1})$ and $p(z_{2})$ do not coincide, so their $U$-orbits are also different. Therefore, if $\dim_{{\mathbb{K}}}\mathfrak{m}^{k}/\mathfrak{m}^{k+1}>1$, there are infinitely many $U$-orbits on $X$, this contradicts our assumption.
It implies that the Hilbert-Samuel sequence has the form $(1,r,1,\ldots,1)$. Now suppose that $r\geq 3$ and consider the map:
$\varphi\colon\mathfrak{m}/\mathfrak{m}^{2}$ $\to\mathfrak{m}^{d}/\mathfrak{m}^{d+1},$
$z+\mathfrak{m}^{2}$ $\mapsto z^{d}+\mathfrak{m}^{d+1}.$
The map $\varphi$ is a morphism between algebraic varieties $\mathfrak{m}/\mathfrak{m}^{2}\simeq{\mathbb{A}}^{r}$ and $\mathfrak{m}^{d}/\mathfrak{m}^{d+1}\simeq{\mathbb{A}}^{1}$. The set $Z:=\varphi^{-1}(0+\mathfrak{m}^{d+1})$ is non-empty, so $\mathrm{dim}(Z)\geq r-1\geq 2$. For all elements $z\in\mathfrak{m}\setminus\{0\}$ with $z+\mathfrak{m}^{2}\in Z$ we have that $p(z)$ is lying in $X$. As previously, when elements $z_{1}+\mathfrak{m}^{2}\in Z$ and $z_{2}+\mathfrak{m}^{2}\in Z$ are not proportional then the $U$-orbits of $p(z_{1})$ and $p(z_{2})$ are different.
Since $\dim(Z)\geq 2$ there are infinitely many $U$-orbits on $X$. This contradicts our assumption, thus $r\leq 2$ and the Hilbert-Samuel sequence of $A$ equals (1,,1)(1$\ldots$,1)( 1 , … , 1 ) or (1,2,1,1)(1,2,1$\ldots$ 1)( 1 , 2 , 1 , … 1 ). ∎
Proposition 6.
Let $(A,U)$ be an $H$-pair and $X$ be the corresponding hypersurface in ${\mathbb{P}}(A).$ Suppose that there are finitely many $U$-orbits in $X$. Then for $n\geq 1$ we have
$A\simeq{\mathbb{K}}[x]/(x^{n+1})\quad\text{or}\quad A\simeq{\mathbb{K}}[x,y]/(xy,x^{3},y^{2}-x^{2}).$
Proof.
First suppose that the Hilbert-Samuel sequence of $A$ equal to (1,1,,1).(1,1$\ldots$,1).( 1 , 1 , … , 1 ) . Then $A$ is generated by one nilpotent element, so $A$ is isomorphic to ${\mathbb{K}}[x]/(x^{n+1}).$
Now consider the case when the Hilbert-Samuel sequence of $A$ is (1,2,,1).(1,2$\ldots$,1).( 1 , 2 , … , 1 ) . Denote by $r$ the maximal number such that $\mathfrak{m}^{r}\neq 0$, where $\mathfrak{m}$ is the maximal ideal in $A$. If $r=1$ then $A\simeq{\mathbb{K}}[x,y]/(x^{2},xy,y^{2})$. In this case $\mathrm{Soc}(A)=\langle x,y\rangle$, this contradicts Corollary 3.
Now consider the case $r>1$. Then there is an element $x\in\mathfrak{m}$ such that $\langle x^{r}\rangle=\mathfrak{m}^{r}$, see [7, Lemma 2.13]. Hence, $\mathfrak{m}=\langle x,x^{2},\ldots,x^{r},y\rangle$ where $y\in\mathfrak{m}\setminus\mathfrak{m}^{2}$ and images of $x$ and $y$ are linearly independent in $\mathfrak{m}/\mathfrak{m}^{2}$. Thus, $xy$ lies in $\mathfrak{m}^{2}$, so $xy=f(x)$, where $f(x)$ is a polynomial divisible by $x^{2}$. We replace $y$ with $y-\frac{f(x)}{x}$ to obtain $xy=0$.
The element $y^{2}$ belongs to $\mathfrak{m}^{2}.$ Thus, $y^{2}=g(x)$, where $g(x)$ is a polynomial divisible by $x^{2}.$ Assume that $y^{2}=0$, then $\mathrm{Soc}(A)=\langle x^{r},y\rangle$, which contradicts Corollary 3. On the other hand, $xy^{2}=(xy)y=0=xg(x).$ It implies that $g(x)=\lambda x^{r}$, where $\lambda\in{\mathbb{K}}\setminus\{0\}.$ We replace $y$ with $\sqrt{\lambda}y$ to get $y^{2}=x^{r}$. Then $A$ is isomorphic to the algebra ${\mathbb{K}}[x,y]/(xy,x^{r+1},y^{2}-x^{r})$.
To complete the proof we should show that $r\leq 2$. Assume the converse, i.e., $r>2$. If we denote by $d\geq 2$ the degree of the hypersurface $X$, then $\mathfrak{m}^{d+1}\subseteq U$. We have
$(y+\alpha x^{2})^{d}=y^{d}+\alpha^{d}x^{2d}\in\mathfrak{m}^{d+1}\quad\text{for all }\alpha\in{\mathbb{K}}.$
Here we use that $y^{2}=x^{r}$ and $y^{3}=0$. Therefore, $p(y+\alpha x^{2})\in X$ for all $\alpha\in{\mathbb{K}}$. The $\mathfrak{m}$-orbit of $(y+\alpha x^{2})$ is the set
$y+\alpha x^{2}+(y+\alpha x^{2})\mathfrak{m}\subseteq y+\alpha x^{2}+\mathfrak{m}^{3}.$
That is, the $\mathfrak{m}$-orbits of the points $p(y+\alpha x^{2})$ do not coincide for different $\alpha$. Hence, if $r>2$ there are infinitely many $U$-orbits on $X$, which leads to a contradiction. ∎
Remark 1.
Note that the algebra ${\mathbb{K}}[x,y]/(xy,x^{3},y^{2}-x^{2})$ is isomorphic to ${\mathbb{K}}[x,y]/(x^{2},y^{2})$. To see this, one should take $\tilde{x}=y-ix$ and $\tilde{y}=y+ix$, then we get
${\mathbb{K}}[x,y]/(xy,x^{3},y^{2}-x^{2})={\mathbb{K}}[\tilde{x},\tilde{y}]/(\tilde{x}^{2},\tilde{y}^{2}).$
We are ready to state our first main result.
Theorem 3.
[…]
$({\mathbb{K}}[x]/(x^{n+1}),\ U_{i}),$ $\text{where }\ U_{i}:=\langle x,x^{2},\ldots,x^{i-1},x^{i+1},\ldots,x^{n}\rangle\ \text{ with }n-1\leq i\leq n,\quad\text{or}$
$({\mathbb{K}}[x,y]/(x^{2},y^{2}),\ W),$ $\text{where }\ W=\langle x,y\rangle.$
To prove Theorem 3, we need the following lemma.
Lemma 1.
  1. (1)
    Let $(A,U)$ be an $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\ U_{i}),$ where
    $U_{i}=\langle x,x^{2}\ldots,x^{i-1},x^{i+1},\ldots,x^{n}\rangle$
    with $i>1$. Consider the corresponding hypersurface $X$. Then there are finitely many $U$-orbits on $X$ if and only if $n-1\leq i\leq n$.
  2. (2)
    Let $(A,U)$ be an $H$-pair $({\mathbb{K}}[x,y]/(x^{2},y^{2}),\ \langle x,y\rangle)$ and $X$ is the corresponding hypersurface. Then there are finitely many $U$-orbits on $X$.
Proof.
[…]
$\{\,p(z)\mid z\in\mathfrak{m}\text{ such that }z^{i}\in U\,\}=p(\mathfrak{m}^{2}).$
By Corollary 1, for each point $p(z)$ from this set the $\mathfrak{m}$-orbit of $p(z)$ is contained in $X$. The total number of $\mathfrak{m}$-orbits on ${\mathbb{P}}(A)$ is finite, see [19, Proposition 3.7]. Each $\mathfrak{m}$-orbit either coincides with an $U$-orbit or is the union of infinite number of $U$-orbits. Therefore, the total number of $U$-orbits in $X$ is finite if and only if for all $z\in\mathfrak{m}^{2}\setminus\{0\}$ the $\mathfrak{m}$-orbit of $p(z)$ is equal to $U$-orbit of $p(z)$. By Proposition 4 this is equivalent to
$\mathrm{Ann}(z)+U=\mathfrak{m},\ \forall z\in\mathfrak{m}^{2}.$
For $z\in\mathfrak{m}^{2}\setminus\mathfrak{m}^{3}$ we have $\mathrm{Ann}(z)=\mathfrak{m}^{n-1}=(x^{n-1})$ and $\mathrm{Ann}(z)\supseteq(x^{n-1})$ for all other $z\in\mathfrak{m}^{2}$. Therefore, the total number of $U$-orbits in $X$ is finite if and only if
$(x^{n-1})+U_{i}=\mathfrak{m}.$
It implies that $i=n$ or $n-1$.
In the case when $(A,U)=({\mathbb{K}}[x,y]/(x^{2},y^{2}),\ \langle x,y\rangle)$, the degree of $X$ is 2. Three $\mathfrak{m}$-orbits are contained in the complement to the open $U$-orbit in $X$. They are $p(x+{\mathbb{K}}xy),p(y+{\mathbb{K}}xy)$ and $p(xy)$. It is easy to see that all these $\mathfrak{m}$-orbits coincide with $U$-orbits. ∎
Proof of Theorem 3.
Let $(A,U)$ be an $H$-pair and suppose that corresponding hypersurface $X\subseteq{\mathbb{P}}^{n}$ contains only a finite number of $U$-orbits. By Proposition 6 and Remark 1 the algebra $A$ is isomorphic to ${\mathbb{K}}[x]/(x^{n+1})$ or ${\mathbb{K}}[x,y]/(x^{2},y^{2})$.
[…]
$U=\langle x+\alpha_{1}x^{n},x^{2}+\alpha_{2}x^{n},\ldots,x^{n-1}+\alpha_{n-1}x^{n}\rangle$
for some $\alpha_{1},\ldots,\alpha_{n-1}\in{\mathbb{K}}$. For all $\beta_{2},\ldots,\beta_{n}\in{\mathbb{K}}$ we consider an automorphism $\varphi$ of $A$, $\varphi\colon x\mapsto x+\beta_{2}x^{2}+\ldots+\beta_{n}x^{n}$. Then
$\varphi(x^{k}+\alpha_{k}x^{n})=(k\beta_{n-k+1}+h_{k}(\beta_{2},\ldots,\beta_{n-k})+\alpha_{k})x^{n}+s_{k}(x),$
where $h_{k}$ and $s_{k}$ are polynomials and the degree of $s_{k}$ is less than $n$.
We take $\beta_{n-k+1}=-\frac{1}{k}(\alpha_{k}+h_{k}(\beta_{2},\ldots,\beta_{n-k}))$ for all $k=1,\ldots,n-1$. Then
$\varphi(x^{k}+\alpha_{k}x^{n})\in\langle x,\ldots,x^{n-1}\rangle\quad\forall k=1,\ldots,n-1.$
Therefore, $\varphi(U)=\langle x,\ldots,x^{n-1}\rangle$.
[…]
$\pi(U)=\langle x+\langle x^{n}\rangle,x^{2}+\langle x^{n}\rangle,\ldots,x^{i-1}+\langle x^{n}\rangle,x^{i+1}+\langle x^{n}\rangle,\ldots\rangle$
for some $i\geq 2$. But then $U=U_{i}$.
Now we consider the case $A\simeq{\mathbb{K}}[x,y]/(x^{2},y^{2})$. If a 2-dimensional subspace $W$ in $\langle x,y,xy\rangle$ generates $A$ then $W=\langle x+\alpha xy,\ y+\beta xy\rangle$. Applying the automorphism of $A$
$x\mapsto x-\alpha xy,\quad y\mapsto y-\beta xy,$
we obtain that $W=\langle x,y\rangle$. Then Lemma 1 completes the proof. ∎
By an $H$-pair we can find the equation of the corresponding hypersurface $X$. For example, we consider the $H$-pair $(A,U)=({\mathbb{K}}[x]/(x^{3}),\ \langle x\rangle)$. Then we apply Theorem 2. If we choose a basis $1,x,x^{2}$ in $A$ then the map $\pi\colon A\to A/U$ can be given as follows:
$z_{0}+z_{1}x+z_{2}x^{2}\mapsto z_{0}+z_{2}x^{2}.$
In this case, the degree of $X$ is 2. If we denote $z=z_{1}x+z_{2}x^{2}$ we obtain
$\ln(1+\frac{z}{z_{0}})=\frac{z}{z_{0}}-\frac{z^{2}}{2z_{0}^{2}}=\frac{z_{1}}{z_{0}}x+\frac{2z_{0}z_{2}-z_{1}^{2}}{2z_{0}^{2}}x^{2}.$
The hypersurface $X$ is then given by the following equation:
$z_{0}^{2}\cdot\pi(\ln(1+\frac{z}{z_{0}}))=z_{0}z_{2}-\frac{1}{2}z_{1}^{2}=0.$
This is a non-degenerate quadric of rank 3. Below we recall the definition of a non-degenerate hypersurface.
Definition 2.
[7, Definition 2.22] Suppose a projective hypersurface $X\subseteq{\mathbb{P}}^{n}$ of degree $d$ is given by an equation $f(z_{0},z_{1},\ldots,z_{n})=0$. Then $X$ is called non-degenerate if there is no linear transformation of variables $z_{0},\ldots,z_{n}$ that reduces the number of variables in $f$ to less than $n+1$.
[…]
Corollary 4.
Let $X\subseteq{\mathbb{P}}^{n}$ be a projective hypersurface admitting an induced additive action with a finite number of orbits.
  1. (1)
    When $n=2$, $X$ is $\mathbb{P}^{1}$ embedded to ${\mathbb{P}}^{2}$ via Veronese embedding of degree 2.
  2. (2)
    When $n=3$, $X$ is one of the following projective surfaces:
    • $(a)$
      ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ embedded to ${\mathbb{P}}^{3}$ as a non-degenerate quadric of rank 4 via Segre embedding;
    • $(b)$
      The non-degenerate cubic $z_{0}^{2}z_{3}-z_{0}z_{1}z_{2}+\frac{z_{1}^{3}}{3}=0$.
    • $(c)$
      The degenerate quadric of rank 3.
  3. (3)
    When $n>3$, $X$ is either a non-degenerate hypersurface $X_{n}$ of degree $n$ or a degenerate hypersurface $Y_{n}$ of degree $n-1$. Moreover, $Y_{n}$ is a projective cone over $X_{n-1}.$
In Table 1 one can find the equations of the resulting hypersurfaces of dimensions 1–4.
$\dim\ X$  The equation of a hypersurface
1    $z_{0}z_{2}-\frac{1}{2}z_{1}^{2}=0$
2    $z_{0}z_{3}-z_{1}z_{2}=0$
2    $z_{0}^{2}z_{3}-z_{0}z_{1}z_{2}+\frac{z_{1}^{3}}{3}=0$
2    $z_{0}z_{2}-\frac{1}{2}z_{1}^{2}=0$
3    $z_{0}^{3}z_{4}-z_{0}^{2}z_{1}z_{3}+\frac{z_{0}^{2}z_{2}^{2}}{2}+z_{0}z_{1}^{2}z_{2}-\frac{z_{1}^{4}}{4}=0$
3    $z_{0}^{2}z_{3}-z_{0}z_{1}z_{2}+\frac{z_{1}^{3}}{3}=0$
4    $z_{0}^{4}z_{5}-z_{0}^{3}z_{1}z_{4}-z_{0}^{3}z_{2}z_{3}+z_{0}^{2}z_{1}^{2}z_{3}+z_{0}^{2}z_{1}z_{2}^{2}-z_{0}z_{1}^{3}z_{2}+\frac{z_{1}^{5}}{5}=0$
4    $z_{0}^{3}z_{4}-z_{0}^{2}z_{1}z_{3}+\frac{z_{0}^{2}z_{2}^{2}}{2}+z_{0}z_{1}^{2}z_{2}-\frac{z_{1}^{4}}{4}=0$
Table 1.
Remark 2.
It was proven in [2] that for each $n\in{\mathbb{N}}$ there is a unique hypersurface in ${\mathbb{P}}^{n}$ of degree $n$ that admits an additive action. It is the hypersurface which corresponds to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\langle x,x^{2},\ldots,x^{n-1}\rangle).$ So we see that except the quadric of rank 4 in ${\mathbb{P}}^{3}$ this is also a unique non-degenerate hypersurface with an additive action with a finite number of orbits.
Remark 3.
Using Theorem 2 it is easy to see that the degenerate hypersurface corresponding to $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\ U)$ with $U_{:}=\langle x,x^{2}\ldots,x^{n-2},x^{n}\rangle$ can be given by the same equation in ${\mathbb{P}}^{n}$ as the non-degenerate hypersurface in ${\mathbb{P}}^{n-1}$ corresponding to $H$-pair $({\mathbb{K}}[x]/(x^{n}),\ U)$ with $U:=\langle x,x^{2}\ldots,x^{n-2}\rangle$.
4. Properties of hypersurfaces
4.1. Orbits
In this section we describe the structure of orbits on hypersurfaces that we found in the previous section.
Let $A$ be a local algebra of dimension $n+1$ and $\mathfrak{m}$ is the maximal ideal in $A$. Then the $\mathfrak{m}$-orbit in ${\mathbb{P}}(A)={\mathbb{P}}^{n}$ of an element $p(z)$ for $z\in A$ is the set
$O_{\mathfrak{m}}(p(z))=\{p(w)\mid w\ \text{is associated with }z\}.$
Here by $O_{\mathfrak{m}}(y)$ for $y\in{\mathbb{P}}^{n}$ we mean the $\mathfrak{m}$-orbit of $y$. Hence, for $A={\mathbb{K}}[x]/(x^{n+1})$ we have $n+1$ $\mathfrak{m}$-orbits:
$O_{\mathfrak{m}}(p(1)),O_{\mathfrak{m}}(p(x)),\ldots,O_{\mathfrak{m}}(p(x^{n})).$
Now we consider the non-degenerate hypersurface $X$ corresponding to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),U)$, where $U=\langle x,x^{2},\ldots x^{n-1}\rangle.$ Since the number of $U$-orbits in $X$ is finite by Proposition 4 all $U$-orbits in $X$ except $U$-orbit of $p(1)$ coincide with $\mathfrak{m}$-orbits.
Among points $p(x),\ldots,p(x^{n})$ exactly points $p(x^{2}),\ldots,p(x^{n})$ belong to $X$. So there are exactly $n$ $U$-orbits in $X$:
$O_{U}(p(1)),O_{U}(p(x^{2})),\ldots,O_{U}(p(x^{n})).$
Similarly, here we denote by $O_{U}(y)$ for $y\in X$ the $U$-orbit of $y$.
For $k\geq 2$ we have
$\mathrm{dim}\ O_{U}(p(x^{k}))=\dim O_{\mathfrak{m}}(p(x^{k}))=\dim\mathfrak{m}-\dim\mathrm{St}_{\mathfrak{m}}(p(x^{k}))=$
$=n-\dim(\mathrm{Ann}(p(x^{k}))=n-\dim\langle x^{n-k+1},\ldots,x^{n}\rangle=n-k.$
At the same time, $O_{U}(p(x^{k}))\subseteq p(\langle x^{k},\ldots,x^{n}\rangle).$ The last set is irreducible, closed in $X$ and has the dimension $n-k$. So $\overline{O_{U}(p(x^{k}))}=p(\langle x^{k},\ldots,x^{n}\rangle).$ It implies that $O_{U}(p(x^{k}))\subseteq\overline{O_{U}(p(x^{l}))}$ if and only if $l\leq k.$
All these arguments also work for the case of the degenerate hypersurface corresponding to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-2},x^{n}\rangle).$ So we obtain the following proposition.
Proposition 7.
Let $X$ be the hypersurface corresponding to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-1}\rangle)$ or $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-2},x^{n}\rangle).$ Then there are exactly $n$ $U$-orbits $O_{0},\ldots,O_{n-1}$ in $X$ with $\dim O_{i}=i$ and $O_{i}\subseteq\overline{O_{j}}$ if and only if $j\geq i$.
It is easy to check that there are 4 orbits in ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}.$ One is two-dimensional and open, two orbits are 1-dimensional and one orbit is a fixed point which is contained in the closures of all others orbits.
4.2. Smoothness and normality
It is clear that ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ is smooth and normal. By [4, Proposition 4] a smooth hypersurface admitting an induced additive action is a non-degenerate quadric. So the only smooth hypersurfaces admitting an induced additive actions with a finite number of orbits are non-degenerate quadrics of rank 3 and 4 in ${\mathbb{P}}^{2}$ and ${\mathbb{P}}^{3}$ respectively.
[…]
$z_{0}^{d}\pi(\ln(1+\frac{z}{z_{0}}))=\sum_{k=1}^{d}z_{0}^{d-k}f_{k}=0,$
where $f_{k}$ is a homogeneous polynomial of degree $k$ and $d$ is the degree of the hypersurface; see Theorem 2. Let $f_{d}=p_{1}^{a_{1}}\ldots p_{r}^{a_{r}}$, where $p_{1},\ldots,p_{r}$ are distinct coprime irreducible polynomials and $a_{i}>0.$ Denote $\tilde{f}_{d}=\frac{f_{d}}{p_{1}\ldots p_{r}}.$
Proposition 8.
[2, Proposition 3] The hypersurface $X$ is normal if and only if the polynomials $\tilde{f}_{d}$ and $f_{d-1}$ are coprime.
It implies that a hypersurface $\{F=0\}\subseteq{\mathbb{P}}^{n}$ admitting an induced additive action is normal if and only if the hypersurface $\{F=0\}\subseteq{\mathbb{P}}^{n+1}$ is normal. So it is enough to check normality for the hypersurfaces corresponding to $H$-pairs $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-1}\rangle)$. In this case $d=n$ and we have
$\tilde{f_{d}}=(-1)^{n}\frac{z_{1}^{n-1}}{n},\ \text{and}\ f_{d-1}=(-1)^{n-1}z_{1}^{n-2}z_{2}.$
Here $z_{0},z_{1}\ldots,z_{n}$ are coordinate functions on ${\mathbb{P}}^{n}$ corresponding to the basis $1,x,\ldots,x^{n}$ in ${\mathbb{K}}[x]/(x^{n+1}).$ These polynomials are coprime if and only if $n=2.$ So we obtain the following result.
Proposition 9.
  1. (1)
    Let $X\subseteq{\mathbb{P}}^{n}$ be a smooth hypersurface admitting an additive action with a finite number of orbits. Then either $n=2$ and $X=\{z_{2}z_{0}-\frac{1}{2}z_{1}^{2}=0\}$ or $n=3$ and $X=\{z_{0}z_{3}-z_{1}z_{2}=0\}.$
  2. (2)
    Let $X\subseteq{\mathbb{P}}^{n}$ be a normal hypersurface admitting an additive action with a finite number of orbits. Then either $n=2$ and $X=\{z_{2}z_{0}-\frac{1}{2}z_{1}^{2}=0\}$ or $n=3$ and $X=\{z_{2}z_{0}-\frac{1}{2}z_{1}^{2}=0\}$ or $X=\{z_{0}z_{3}-z_{1}z_{2}=0\}$.
4.3. The number of additive actions
By [7, Theorem 2.32] a non-degenerate hypersurface in ${\mathbb{P}}^{n}$ admits at most one induced additive action. At the same time, it was proven in [10] that if a degenerate hypersurface in ${\mathbb{P}}^{n}$ admits an induced additive action then it admits at least two non-isomorphic induced additive actions. Here we will prove that if a degenerate hypersurface in ${\mathbb{P}}^{n}$ admits an induced additive action with a finite number of orbits then it admits exactly two induced additive actions.
[…]
Let $y$ be a preimage of $x$ in $W$ with respect to the canonical projection $B\to B/P$. Denote by $z$ a basis in $P$. Then $z\in\mathrm{Soc}(B).$
The image of $y^{n}$ in $B/P$ is zero so $y^{n}\in\langle z\rangle.$ If $y^{n}\neq 0$ then up to a scalar $y^{n}=z.$ In this case $B={\mathbb{K}}[y]/(y^{n+1})$ and $W=\langle y,\ldots,y^{n-2},y^{n}\rangle$. If $y^{n}=0$ then $B={\mathbb{K}}[y,z]/(y^{n},z^{2},zy)$ and $W=\langle y,\ldots,y^{n-2},z\rangle$. So we have proved the following proposition.
Proposition 10.
Let $X$ be a hypersurface in ${\mathbb{P}}^{n}$ which corresponds to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-2},x^{n}\rangle)$. Then there are exactly two induced additive actions on $X$. One corresponds to the $H$-pair $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-2},x^{n}\rangle)$ and the other one corresponds to the $H$-pair $({\mathbb{K}}[y,z]/(y^{n},z^{2},zy),\langle y,\ldots,y^{n-2},z\rangle)$.
It is not difficult to describe the additive action corresponding to the $H$-pair $({\mathbb{K}}[y,z]/(y^{n},z^{2},zy),\langle y,\ldots,y^{n-2},z\rangle)$. Let $X_{n-1}$ be the hypersurface in ${\mathbb{P}}^{n-1}$ corresponding to the $H$-pair $({\mathbb{K}}[y]/(y^{n}),\langle y,y^{2},\ldots,y^{n-2}\rangle)$. We denote by $z_{0},\ldots,z_{n-1}$ the homogeneous coordinates on ${\mathbb{P}}^{n-1}$ corresponding to the basis $1,y,\ldots,y^{n-1}$. Then the element
$(s_{1},\ldots,s_{n-2})=s_{1}y+s_{2}y^{2}+\ldots+s_{n-2}y^{n-2}\in U$
acts on $X_{n-1}$ in the following way
$(s_{1},\ldots,s_{n-2})\circ[z_{0}:\ldots:z_{n-1}]=[z_{0}:z_{1}+g_{1}:\ldots:z_{n-1}+g_{n-1}],$
where $g_{1},\ldots,g_{n-1}$ are polynomials in variables $z_{0},\ldots,z_{n},s_{1},\ldots,s_{n-1}.$ This action is linear, so the polynomials $g_{1},\ldots,g_{n-1}$ are homogenous polynomials of degree 1 in variables $z_{0},\ldots,z_{n-1}$. Moreover, one can check that the polynomial $g_{i}$ does not depend on $z_{i},\ldots,z_{n}.$
Suppose that $X_{n-1}$ is given in ${\mathbb{P}}^{n-1}$ by a homogeneous polynomial $F$. Then the hypersurface $Y_{n}$ corresponding to the $H$-pair $({\mathbb{K}}[y,z]/(y^{n},z^{2},zy),\langle y,\ldots,y^{n-2},z\rangle)$ is given in ${\mathbb{P}}^{n}$ by the same polynomial $F$ and the corresponding additive action is given by the following formula:
$(s_{1},\ldots,s_{n-2},s_{n-1})\circ[z_{0}:\ldots:z_{n-1}:z_{n}]=[z_{0}:z_{1}+g_{1}:\ldots:z_{n-1}+g_{n-1}:z_{n}+s_{n-1}z_{0}],$
where $z_{0},\ldots,z_{n}$ are homogeneous coordinates on ${\mathbb{P}}^{n}$ corresponding to the basis $1,y,y^{2},\ldots,y^{n-1},z$ of ${\mathbb{K}}[y,z]/(y^{n},z^{2},zy)$ and
$(s_{1},\ldots,s_{n-1})=s_{1}y+s_{2}y^{2}+\ldots+s_{n-2}y^{n-2}+s_{n-1}z\in U.$
We denote by $O_{i}$ the orbit in $X_{n-1}$ of dimension $i$. Then the open orbit $O_{n-2}$ in $X_{n-1}$ is the set
$O_{n-2}=\{[z_{0}:\ldots:z_{n-1}]\in{\mathbb{P}}^{n-1}\mid F(z_{0},\ldots,z_{n-1})=0\ \text{and}\ z_{0}\neq 0\}.$
Then the open orbit $O$ in $Y_{n}$ is the set
$O=\{[z_{0}:\ldots:z_{n-1}:z_{n}]\in{\mathbb{P}}^{n-1}\mid F(z_{0},\ldots,z_{n-1})=0\ \text{and}\ z_{0}\neq 0\}.$
When $i<n-2$ the orbit $O_{i}$ in $X_{n-1}$ is the orbit of a point $P_{i}=[0:0:\ldots 1:0:\ldots 0]$ where 1 stands at the coordinate $z_{n-i-1}.$ There are infinitely many orbits of dimension $i$ in $Y_{n}$. They are orbits of the points
$P_{i,c}=[0:0:\ldots 0:1:0:\ldots 0:c]$
(again, 1 stands at the coordinate $z_{n-i-1}).$ If the closure $\overline{O_{i}}$ in $X_{n-1}$ is given by the set of homogeneous polynomials
$F_{i,1},\ldots,F_{i,r(i)}$
then the closure of the orbit of $P_{i,c}$ in $Y_{n}$ is given by the set of polynomials
$F_{i,1},\ldots,F_{i,r(i)},z_{n}-cz_{n-i-1}.$
When $i>0$ the closure of the orbit of $P_{i,c}$ contains the orbit of $P_{i,c}$ and the orbits $O_{0},\ldots,O_{i-1}.$ Here, we assume that $X_{n-1}$ is the subset of $Y_{n}$ which is given by the equation $z_{n}=0.$
4.4. Limit points of one-parameter subgroups
Let $X$ be a complete variety and $\alpha:\mathbb{G}_{a}^{n}\times X\to X$ be an additive action on $X$. Let $O$ be the open orbit. We say that the additive action $\alpha$ satisfies OP-condition (one-parameter subgroups condition) if for every point $x\in X$ there is a point $y\in O$ and a one-dimensional subgroup $S\subseteq\mathbb{G}_{a}^{n}$ such that $x\in\overline{Sy}.$
In [13] normal complete varieties with an additive action with a finite number of orbits and OP-condition were described. More precisely, the following holds.
Theorem 4.
[13, Theorem A] Let $X$ be a complete variety and $\alpha:\mathbb{G}_{a}^{n}\times X\to X$ is an additive action. Then the following conditions are equivalent.
  1. (1)
    There are finitely many orbits on $X$ with respect to $\alpha$ and $\alpha$ satisfies OP-condition.
  2. (2)
    The variety $X$ is a matroid Schubert variety and $\alpha$ is the corresponding additive action on $X$.
One can find the definition of a matroid Schubert variety in the introduction of[13]. In this section we check what additive actions on projective hypersurfaces with a finite number of orbits satisfy OP-condition.
Suppose $X$ is a complete variety and $\alpha:\mathbb{G}_{a}^{n}\times X\to X$ is an effective additive action on $X$ with an open orbit $O$. Then $O\simeq\mathbb{G}_{a}^{n}$ and we can assume that the group $\mathbb{G}_{a}^{n}$ is a subset in $X$. Then $\alpha$ satisfies OP-condition if and only if for every $\mathbb{G}_{a}^{n}$-orbit $O^{\prime}$ there is a one-dimensional subgroup $S\subseteq\mathbb{G}_{a}^{n}$ such that $\overline{S}\cap O^{\prime}\neq\emptyset$.
Now we assume that $X$ is the projective hypersurface corresponding to the $H$-pair $(A,U)=({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-1}\rangle)$ and $\alpha$ is the respective additive action. Consider a one-dimensional subgroup $S=\langle\alpha_{1}x+\ldots+\alpha_{n-1}x^{n-1}\rangle\subseteq U.$ The image of $S$ in $X$ is the subset $p(\mathrm{exp}(S)).$ Let $z_{0},\ldots,z_{n}$ be the coordinates in ${\mathbb{P}}(A)={\mathbb{P}}^{n}$ corresponding to the basis $1,x,\ldots,x^{n}$ of $A$. We have
$\mathrm{exp}(S)=\{1+t(\alpha_{1}x+\ldots+\alpha_{n-1}x^{n-1})+\ldots+\frac{t^{n}\alpha_{1}^{n}x^{n}}{n!}\mid t\in{\mathbb{K}}\}$
and $p(\mathrm{exp}(S))$ is a set of the form $\{[g_{0}(t):\ldots g_{n}(t)]\mid t\in{\mathbb{K}}\}$, where $g_{i}$ is a polynomial of degree $i$.
If we consider a homogeneous polynomial $F(z_{0},\ldots,z_{n})$ which is equal to zero on $\overline{\exp(S)}$ then the polynomial $f(t)=F(g_{0}(t),\ldots,g_{n}(t))$ is zero. It implies that there are no monomials of the form $z_{n}^{k}$ in $F$. So $F(0,\ldots,0,1)=0$. Therefore, $[0:\ldots:1]\in\overline{\exp(S)}.$
The set $\overline{\exp(S)}$ is the union of $\exp(S)$ and [0::0:1][0$\ldots$:0:1][ 0 : … : 0 : 1 ]. Indeed, the group $S\simeq{\mathbb{G}}_{a}$ acts on $\overline{\exp(S)}$ and $\exp(S)$ is an open orbit in $\overline{\exp(S)}$. So $\overline{\exp(S)}\setminus\exp(S)$ is a finite set of fixed points. But by [12, Corollary 1] the set of ${\mathbb{G}}_{a}$-fixed points on a complete variety is connected. So there is only one point in $\overline{\exp(S)}\setminus\exp(S).$
Therefore, for any one-dimensional subgroup $S$ in $U$ the set $\overline{\exp(S)}$ intersects only two $U$-orbits: the open orbit and the fixed point $p(x^{n}).$ Hence, OP-condition holds only when $n=2.$ The same arguments show that OP-condition never holds for the $H$-pairs $({\mathbb{K}}[x]/(x^{n+1}),\langle x,\ldots,x^{n-2},x^{n}\rangle)$ and it is easy to check that OP-condition holds for the additive action on ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$. Thus, we obtain the following proposition.
Proposition 11.
Let $X$ be a projective hypersurface which admits an additive action with a finite number of orbits and satisfying OP-condition. Then $X$ is either $\{z_{2}z_{0}-\frac{1}{2}z_{1}^{2}=0\}\subseteq{\mathbb{P}}^{2}$ or $\{z_{0}z_{3}-z_{1}z_{2}=0\}\subseteq{\mathbb{P}}^{3}.$
Since ${\mathbb{P}}^{1}$ and ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$ are normal varieties, we obtain the following corollary.
Corollary 5.
Let $X$ be a matroid Schubert variety which can be embedded as a hypersurface in a projective space. Then $X$ is ${\mathbb{P}}^{1}$ or ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$.